You get an electric shock in which 9.0 x 10-12 Coulombs is transferred. How many electrons were involved in that "shock" you received?
Answer #1

electric charge Q = 9.0 x 10-12 Coulombs

number of electrons n = Q/charge of electron

n = Q/e = 9.0 x 10-12/1.6x10-19

n = 5.625*107 electrons